# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

### v3

@@ II.1 Localized phases
We seek to construct a continuous path of locality preserving unitaries $V_{t}$ such that $V_{t}$ conjugates the LIOMs of $H_{0}$ into LIOMs of $H_{t}$. A unitary $V$ is called locality preserving if $VAV^{\dagger}$ is localized near the support of $A$ for all local operators $A$. In a physical context, this means $VAV^{\dagger}$ has some operator decomposition where the norm of a term in the decomposition decays with its distance from the support of $A$. Any finite time evolution under a local Hamiltonian is locality preserving in this sense. Mathematically, it is convenient to deal with strictly locality preserving unitaries, where the support of $VAV^{\dagger}$ grows by a finite distance $r$. This includes all finite depth local quantum circuits. Locality-preserving unitaries are also called quantum cellular automata (QCA), and we will use this terminology in the rest of the paper.
@@ II.3 Driven phases
Most examples of ALT phases in the literature are driven systems [24, 23, 49, 25, 29, 50, 31]. We specialize to the case of multi-tone time dependence [103, 28, 104] (which includes Floquet driving as a special case), such that the Hamiltonian can be parameterized as
@@
Here, $U(t)$ is the time evolution operator, the Floquet Hamiltonian $H_{F}=\sum_{j}h_{j}\sigma^{z}_{j}+\cdots$ is trivially MBL, and the micromotion operator $V({\vec{\theta}})$ is a family of QCA parameterized by $\vec{\theta}$}. Thus, in complete analogy with the static case, we have a correspondence between multi-tone-driven localized phases and parameterized families of QCA, $V({\vec{\theta}})$, up to continuous deformations. That is, homotopy classes of continuous maps from the $n$-dimensional torus to $\mathrm{QCA}_{d}$. This reduces to the $n$th homotopy group $\pi_{n}(\mathrm{QCA}_{d})$ when ignoring weak invariants [105, 106] corresponding to low dimensional cross sections of the torus. Collapsing some of these cross-sections to a point reduces the torus to a sphere, and the $n$th homotopy group precisely corresponds to families $V({\vec{\theta}})$ parameterized by a variable on a sphere, up to continuous deformations.
@@
As in the case of static systems, the parameterization by QCA is not unique, so that the classification of ALT phases is properly a quotient of $\ker e^{*}_{0}$. Further details are provided in Appendix A.
@@ III.1 Loops and $\Omega$-spectra
The relationship Eq. (6) is reminiscent of a widely held conjecture that invertible states form an $\Omega$-spectrum [101, 100, 63, 64, 65]. In topology, an $\Omega$-spectrum refers to a sequence of topological spaces $(Y_{d})_{d\in{\mathbb{Z}}}$ together with (weak) homotopy equivalences (that is, continuous maps which induce isomorphisms on all homotopy groups)
@@
The hypothesis that invertible states form an $\Omega$-spectrum implies immediately that $\pi_{n}(\mathrm{inv}_{d})=\pi_{n-1}(\mathrm{inv}_{d-1})$. Further, it is believed that the classification of symmetry-enriched invertible states with symmetry group $G$ are completely classified by the generalized cohomology of the classifying space $BG$ [i.e. $h^{d}_{\mathrm{inv}}(BG)$[100, 63, 65, 64].
@@
Indeed, continuous deformations of the bulk loop $\gamma$ produce continuous deformations of $v$ [107], so the topological invariants of $v$ may also be considered invariants for $\gamma$. The difficult steps needed to prove Eq. (9) are to show that:
@@ III.3 Computing the classification
The connected components $\pi_{0}(\mathrm{QCA}_{d})$ are known (for spin systems) for $d=1$ [58] and $d=2$ [61]. There are sound conjectures for $d=3$ [56, 109, 55]. The space $\mathrm{QCA}_{0}$ is a stabilization of the projective unitary group by stacking ancillae. We recall existing classifications and compute the homotopy groups of $\mathrm{QCA}_{0}$ in Sec. III.3.1 and Sec. III.3.2 respectively.
Thus, we have all the homotopy groups $\pi_{n}(\mathrm{QCA}_{d})$ for spin systems with $d-n\leq 2$, and can make a strong conjecture for dn=3d-n=3. The results are listed in Tab. 1, and the corresponding classification of localized phases (excluding permutation QCA) is listed in Tab. 2. Comparing to the classification of invertible states shows that all the nontrivial phases in the tables are ALT phases.
@@ 33 Witt 0 $\log{\mathbb{Q}}^{\times}$ 0 $\mathbb{Q}$$\mathbb{Z}$} 0
Table 2: The classification of $n$-tone driven $d$-dimensional localized phases of spins with short-range entanglement. The classifying group differs from $\pi_{n}(\mathrm{QCA}_{d})$ for $(d,n)\in\{(0,1),(1,0)\}$ due to the non-uniqueness of the parameterization by QCA (Appendix A.3). All nontrivial elements of this table correspond to ALT phases.
@@ III.3.1 Connected components
Each prime Hilbert space dimension is labelled by $\log p$, and the integer coefficients give the shift index for that qudit species.
@@
The classification of three dimensional QCA is still open. It is conjectured that $\pi_{0}(\mathrm{QCA}_{3})$ is isomorphic to the Witt group of modular tensor categories [56, 109, 55, 110]—that is, it has the same group structure as stacking two-dimensional topological orders, modulo an equivalence relation which identifies phases which may share a gapped edge. Presently, Clifford QCA (which map Pauli operators to Pauli operators) have been classified [109], and the classification is consistent with the Witt group conjecture for $\pi_{0}(\mathrm{QCA}_{3})$, but it has not been formally verified that all these phases of Clifford QCA remain nontrivial as non-Clifford QCA. Nonetheless, the existence of nontrivial elements of $\pi_{0}(\mathrm{QCA}_{3})$ is strongly supported by the constructions of Refs. [56, 55]. The models of those references can, indeed, be identified with two-dimensional topologically ordered phases. In fact, the MBL Hamiltonians corresponding to those models exhibit that topological order (specifically, the associated anyonic excitations) on their two dimensional surfaces [48, 56, 55, 110].
@@ III.3.2 Zero dimensions
The space of zero dimensional QCA with finite Hilbert space dimension, say $N$, is naturally identified with the projective unitary group,
@@
There are even integral formulae for winding number invariants indexing the elements of these groups. These formulae involve first making a lift from projective unitaries to linear unitaries. Indeed, given any continuous map $V:S^{n}\to PU(N)$ (where $S^{n}$ is the $n$-dimensional sphere) with $n\neq 2$ it is possible to find a lift $\tilde{V}:S^{n}\to U(N)$ to the usual unitary group. (For $n=2$, we already have that $\pi_{2}(PU(N))\cong 0$, so we don’t need any formula to compute the homotopy class.) That is, if $q:U(N)\to PU(N)$ is the quotient map from the unitary group to the projective unitaries, we have $q\tilde{V}=V$. [^3] In general, the existence of such a lift is guaranteed by $V^{*}(\omega)\in H^{2}[S^{n},{\mathbb{Z}}]$ being trivial, where $\omega\in H^{2}[PU(N),{\mathbb{Z}}]$ is the Euler class for $q:U(N)\to PU(N)$ and $V^{*}$ is the map induced by $V$ on singular cohomology. But $H^{2}[S^{n},{\mathbb{Z}}]\cong 0$ for $n\neq 2$, so the map always exists for our cases of interest. Defining $A=(\mathrm{d}\tilde{V})\tilde{V}^{\dagger}$, where $\mathrm{d}$ is the exterior derivative on $S^{n}$, the winding number invariants are given by [31, 111]
@@
A similar directed system to Eqs. (2223) defines the limit $\mathrm{QCA}_{0}^{\mathcal{R}}$ describing the stabilization of QCA by the ancillae in $\mathcal{R}$. The only difference is that now the system proceeds through increasing $\rho\in\mathcal{R}$, rather than $N\in{\mathbb{N}}$. Thus, not all denominators may occur in the rescaled winding numbers. Defining the algebraic localization (not to be confused with localization in physical space)
@@ IV Symmetry enrichment
where $\mathrm{Rep}(G)$ is the tensor category of unitary equivalence classes of finite dimensional unitary representations of $G$. (This can be replaced with projective representations if appropriate.)
@@
Indeed, the ingredient which allows us to complete the classification of symmetric QCA in one-dimension is to recognize that the elements of $\mathcal{R}$ can be treated analogously to the dimensions of available qudits. We can define a shift index as for one-dimensional QCA without symmetry, only rather than considering ratios of dimensions we deal with formal ratios of representations. The resulting classification is similar to that in Tab. 1, except that rational numbers $p/q$ are replaced by such formal ratios, $\rho/\sigma$. In particular, if ancillae carrying any finite dimensional representation can be added to the system, one-dimensional symmetric QCA are completely classified by the Grothendieck group of projective representations of $G$.
We summarize a few examples. The classification of QCA without symmetry [58] can be regarded as a special case where $G=0$ is the trivial group. In this case, projective representations are just (projective) vector spaces classified by their (nonzero) dimension $N\in{\mathbb{N}}^{\times}$, which multiply under tensor product. The Grothendieck group consists of formal ratios of dimensions, and is just the group of positive rationals with multiplication, $\mathbb{Q}$$^{\times}$. In Tab. 1, we express this as $\log{\mathbb{Q}}^{\times}$ so that the group operation is addition. Number-conserving, that is $U(1)$-symmetric, QCA were found to be classified by rational polynomials in Ref. [26]. This also follows from the observation that projective representations of $U(1)$ are classified by Laurent polynomials with positive integer coefficients, and that the Grothendieck group thereof precisely consists of rational polynomials. Finally, the classification with finite $G$ and all ancillae available collapses to the classification of SPT states, $H^{2}[G;U(1)]$, times the shift index classification, $\log{\mathbb{Q}}^{\times}$ [66, 67]. This occurs because the representation ring of a finite group has zero divisors, which causes many elements to become identified in the Grothendieck group.
@@ IV.1 Loops with symmetry
is satisfied. Indeed, we will show that if the swindle map $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ is in fact a homotopy equivalence, then it also provides the homotopy equivalence for the case with symmetry. Togther with Sec. IV.2 and Appendix C, the $\Omega$-spectrum condition gives us the classification of $\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})$ for $d-n\leq 1$.
@@
which classifies parameterized families of symmetric QCA.
@@ IV.2.1 Structure of one-dimensional symmetric QCA
The product of two consecutive $v_{x}v_{x+1}$ is, by Eq. (48), unitarily equivalent to the representation $\rho_{x}\rho_{x+1}$. However, this only implies that $v_{x}$ and $v_{x+1}$ are projective representations of $G$. The classifications of Ref. [66, 26, 67] rely on the following observation: $V$ is equal to a finite depth circuit with symmetric unitary gates (which implies they are connected by a symmetric path) and the same circuit geometry if and only if there are functions $\phi_{x}:G\to U(1)$ such that $e^{i\phi_{x}}v_{x}$ is unitarily equivalent to $\rho_{x}$ for all $x$. That is, the projective representation $v_{x}$ can be lifted to an ordinary unitary representation, and that representation is precisely $\rho_{x}$. A more concise statement of the condition is that $v_{x}$ and $\rho_{x}$ are isomorphic when both regarded as projective representations.
@@ IV.2.3 Classification
This classifying group can be realized as an extension of a linear part by a projective part. The linear part is a localization of the submonoid of $\mathrm{Rep}_{\mathrm{proj}}(G)$ consisting of projective representations that admit lifts to linear representations. This submonoid is isomorphic to $\mathrm{Rep}(G)/X_{G}$, where $X_{G}$ is the group of units (invertible elements) in $\mathrm{Rep}(G)$—that is, one-dimensional representations—which all become identified as projective representations [67]. The projective part is classified by $H^{2}[G;U(1)]$, as is well known. The induced homomorphism from evaluation on a symmetric product state, $e_{0}^{*}$, just corresponds to sending a projective representation to its classifying cohomology class in $H^{2}[G;U(1)]$, and thus reproduces the group cohomology classification of invertible states in one dimension [71, 66].
@@ V Models
In Sec. V.1 we briefly describe a simple class of examples, building on the anomalous Floquet phases of Refs. [22, 23]. These examples are built from swindles applied to a translation QCA, and give a simple picture illustrating the anomalous edge dynamics of ALT phases. Sec. V.2 constructs a more experimentally viable implementation of the swindle map, and Sec. V.3 applies it to a very detailed investigation of a one-dimensional SALT phase with two drives and $U(1)$-symmetry, the quasiperiodic energy pump (QP pump) [31, 32, 53]. We include explicit expressions for the Hamiltonian and quantized observables which reveal the anomalous edge dynamics. The model we obtain through the swindle construction is relatively simple, at least in its operator content, paving the way for experiments on this phase.

### v4

@@ II.1 Localized phases
We seek to construct a continuous path of locality preserving unitaries $V_{t}$ such that $V_{t}$ conjugates the LIOMs of $H_{0}$ into LIOMs of $H_{t}$. A unitary $V$ is called locality preserving if $VAV^{\dagger}$ is localized near the support of $A$ for all local operators $A$. In a physical context, this means $VAV^{\dagger}$ has some operator decomposition where the norm of a term in the decomposition decays with its distance from the support of $A$. Any finite-time evolution under a local Hamiltonian is locality preserving in this sense. Mathematically, it is convenient to deal with strictly locality preserving unitaries, where the support of $VAV^{\dagger}$ grows by a finite distance $r$. This includes all finite depth local quantum circuits. Locality-preserving unitaries are also called quantum cellular automata (QCA), and we will use this terminology in the rest of the paper.
@@ II.3 Driven phases
Most examples of ALT phases in the literature are driven systems [24, 23, 49, 25, 29, 50, 31]. We specialize to the case of multi-tone time dependence [103, 28, 104] (which includes Floquet driving as a special case), such that the Hamiltonian can be parametrized as
@@
Here, $U(t)$ is the time evolution operator, the Floquet Hamiltonian $H_{F}=\sum_{j}h_{j}\sigma^{z}_{j}+\cdots$ is trivially MBL, and the micromotion operator $V({\vec{\theta}})$ is a family of QCA parametrized by $\vec{\theta}$}. Thus, in complete analogy with the static case, we have a correspondence between multi-tone-driven localized phases and parametrized families of QCA, $V({\vec{\theta}})$, up to continuous deformations. That is, homotopy classes of continuous maps from the $n$-dimensional torus to $\mathrm{QCA}_{d}$. This reduces to the $n$th homotopy group $\pi_{n}(\mathrm{QCA}_{d})$ when ignoring weak invariants [105, 106] corresponding to low dimensional cross sections of the torus. Collapsing some of these cross-sections to a point reduces the torus to a sphere, and the $n$th homotopy group precisely corresponds to families $V({\vec{\theta}})$ parametrized by a variable on a sphere, up to continuous deformations.
@@
As in the case of static systems, the parametrization by QCA is not unique, so that the classification of ALT phases is properly a quotient of $\ker e^{*}_{0}$. Further details are provided in Appendix A.
@@ III.1 Loops and $\Omega$-spectra
The relationship (6) is reminiscent of a widely held conjecture that invertible states form an $\Omega$-spectrum [101, 100, 63, 64, 65]. In topology, an $\Omega$-spectrum refers to a sequence of topological spaces $(Y_{d})_{d\in{\mathbb{Z}}}$ together with (weak) homotopy equivalences (that is, continuous maps which induce isomorphisms on all homotopy groups)
@@
The hypothesis that invertible states form an $\Omega$-spectrum implies immediately that $\pi_{n}(\mathrm{inv}_{d})=\pi_{n-1}(\mathrm{inv}_{d-1})$. Further, it is believed that the classification of symmetry-enriched invertible states with symmetry group $G$ are completely classified by the generalized cohomology of the classifying space $BG$ [i.e., $h^{d}_{\mathrm{inv}}(BG)$[100, 63, 65, 64].
@@
Indeed, continuous deformations of the bulk loop $\gamma$ produce continuous deformations of $v$ [107], so the topological invariants of $v$ may also be considered invariants for $\gamma$. The difficult steps needed to prove Eq. (9) are to show the following:
@@ III.3 Computing the classification
The connected components $\pi_{0}(\mathrm{QCA}_{d})$ are known (for spin systems) for $d=1$ [58] and $d=2$ [61]. There are sound conjectures for $d=3$ [56, 109, 55]. The space $\mathrm{QCA}_{0}$ is a stabilization of the projective unitary group by stacking ancillae. We recall existing classifications and compute the homotopy groups of $\mathrm{QCA}_{0}$ in Secs. III.3.1 and III.3.2 respectively.
Thus, we have all the homotopy groups $\pi_{n}(\mathrm{QCA}_{d})$ for spin systems with $d-n\leq 2$, and can make a strong conjecture for dn=3d-n=3. The results are listed in Table 1, and the corresponding classification of localized phases (excluding permutation QCA) is listed in Table 2. Comparing to the classification of invertible states shows that all the nontrivial phases in the tables are ALT phases.
@@ 33 Witt 0 $\log{\mathbb{Q}}^{\times}$ 0 $\mathbb{Q}$$\mathbb{Z}$} 0
Table 2: The classification of $n$-tone driven $d$-dimensional localized phases of spins with short-range entanglement. The classifying group differs from $\pi_{n}(\mathrm{QCA}_{d})$ for $(d,n)\in\{(0,1),(1,0)\}$ due to the non-uniqueness of the parametrization by QCA (Appendix A.3). All nontrivial elements of this table correspond to ALT phases.
@@ III.3.1 Connected components
Each prime Hilbert space dimension is labeled by $\log p$, and the integer coefficients give the shift index for that qudit species.
@@
The classification of three-dimensional QCA is still open. It is conjectured that $\pi_{0}(\mathrm{QCA}_{3})$ is isomorphic to the Witt group of modular tensor categories [56, 109, 55, 110]—that is, it has the same group structure as stacking two-dimensional topological orders, modulo an equivalence relation which identifies phases which may share a gapped edge. Presently, Clifford QCA (which map Pauli operators to Pauli operators) have been classified [109], and the classification is consistent with the Witt group conjecture for $\pi_{0}(\mathrm{QCA}_{3})$, but it has not been formally verified that all these phases of Clifford QCA remain nontrivial as non-Clifford QCA. Nonetheless, the existence of nontrivial elements of $\pi_{0}(\mathrm{QCA}_{3})$ is strongly supported by the constructions of Refs. [56, 55]. The models of those references can, indeed, be identified with two-dimensional topologically ordered phases. In fact, the MBL Hamiltonians corresponding to those models exhibit that topological order (specifically, the associated anyonic excitations) on their two dimensional surfaces [48, 56, 55, 110].
@@ III.3.2 Zero dimensions
The space of zero-dimensional QCA with finite Hilbert space dimension, say $N$, is naturally identified with the projective unitary group,
@@
There are even integral formulas for winding number invariants indexing the elements of these groups. These formulas involve first making a lift from projective unitaries to linear unitaries. Indeed, given any continuous map $V:S^{n}\to PU(N)$ (where $S^{n}$ is the $n$-dimensional sphere) with $n\neq 2$ it is possible to find a lift $\tilde{V}:S^{n}\to U(N)$ to the usual unitary group. (For $n=2$, we already have that $\pi_{2}(PU(N))\cong 0$, so we do not need any formula to compute the homotopy class.) That is, if $q:U(N)\to PU(N)$ is the quotient map from the unitary group to the projective unitaries, we have $q\tilde{V}=V$. [^3] In general, the existence of such a lift is guaranteed by $V^{*}(\omega)\in H^{2}[S^{n},{\mathbb{Z}}]$ being trivial, where $\omega\in H^{2}[PU(N),{\mathbb{Z}}]$ is the Euler class for $q:U(N)\to PU(N)$ and $V^{*}$ is the map induced by $V$ on singular cohomology. But $H^{2}[S^{n},{\mathbb{Z}}]\cong 0$ for $n\neq 2$, so the map always exists for our cases of interest. Defining $A=(\mathrm{d}\tilde{V})\tilde{V}^{\dagger}$, where $\mathrm{d}$ is the exterior derivative on $S^{n}$, the winding number invariants are given by [31, 111]
@@
A similar directed system to Eqs. (22) and (23) defines the limit $\mathrm{QCA}_{0}^{\mathcal{R}}$ describing the stabilization of QCA by the ancillae in $\mathcal{R}$. The only difference is that now the system proceeds through increasing $\rho\in\mathcal{R}$, rather than $N\in{\mathbb{N}}$. Thus, not all denominators may occur in the rescaled winding numbers. Defining the algebraic localization (not to be confused with localization in physical space)
@@ IV Symmetry enrichment
where $\mathrm{Rep}(G)$ is the tensor category of unitary equivalence classes of finite-dimensional unitary representations of $G$. (This can be replaced with projective representations if appropriate.)
@@
Indeed, the ingredient which allows us to complete the classification of symmetric QCA in one-dimension is to recognize that the elements of $\mathcal{R}$ can be treated analogously to the dimensions of available qudits. We can define a shift index as for one-dimensional QCA without symmetry, only rather than considering ratios of dimensions we deal with formal ratios of representations. The resulting classification is similar to that in Table 1, except that rational numbers $p/q$ are replaced by such formal ratios, $\rho/\sigma$. In particular, if ancillae carrying any finite dimensional representation can be added to the system, one-dimensional symmetric QCA are completely classified by the Grothendieck group of projective representations of $G$.
We summarize a few examples. The classification of QCA without symmetry [58] can be regarded as a special case where $G=0$ is the trivial group. In this case, projective representations are just (projective) vector spaces classified by their (nonzero) dimension $N\in{\mathbb{N}}^{\times}$, which multiply under tensor product. The Grothendieck group consists of formal ratios of dimensions, and is just the group of positive rationals with multiplication, $\mathbb{Q}$$^{\times}$. In Table 1, we express this as $\log{\mathbb{Q}}^{\times}$ so that the group operation is addition. Number-conserving, that is $U(1)$-symmetric, QCA were found to be classified by rational polynomials in Ref. [26]. This also follows from the observation that projective representations of $U(1)$ are classified by Laurent polynomials with positive integer coefficients, and that the Grothendieck group thereof precisely consists of rational polynomials. Finally, the classification with finite $G$ and all ancillae available collapses to the classification of SPT states, $H^{2}[G;U(1)]$, times the shift index classification, $\log{\mathbb{Q}}^{\times}$ [66, 67]. This occurs because the representation ring of a finite group has zero divisors, which causes many elements to become identified in the Grothendieck group.
@@ IV.1 Loops with symmetry
is satisfied. Indeed, we will show that if the swindle map $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ is in fact a homotopy equivalence, then it also provides the homotopy equivalence for the case with symmetry. Together with Sec. IV.2 and Appendix C, the $\Omega$-spectrum condition gives us the classification of $\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})$ for $d-n\leq 1$.
@@
which classifies parametrized families of symmetric QCA.
@@ IV.2.1 Structure of one-dimensional symmetric QCA
The product of two consecutive $v_{x}v_{x+1}$ is, by Eq. (48), unitarily equivalent to the representation $\rho_{x}\rho_{x+1}$. However, this only implies that $v_{x}$ and $v_{x+1}$ are projective representations of $G$. The classifications of Refs. [66, 26, 67] rely on the following observation: $V$ is equal to a finite depth circuit with symmetric unitary gates (which implies they are connected by a symmetric path) and the same circuit geometry if and only if there are functions $\phi_{x}:G\to U(1)$ such that $e^{i\phi_{x}}v_{x}$ is unitarily equivalent to $\rho_{x}$ for all $x$. That is, the projective representation $v_{x}$ can be lifted to an ordinary unitary representation, and that representation is precisely $\rho_{x}$. A more concise statement of the condition is that $v_{x}$ and $\rho_{x}$ are isomorphic when both are regarded as projective representations.
@@ IV.2.3 Classification
This classifying group can be realized as an extension of a projective part by a linear part. The linear part is a localization of the submonoid of $\mathrm{Rep}_{\mathrm{proj}}(G)$ consisting of projective representations that admit lifts to linear representations. This submonoid is isomorphic to $\mathrm{Rep}(G)/X_{G}$, where $X_{G}$ is the group of units (invertible elements) in $\mathrm{Rep}(G)$—that is, one-dimensional representations—which all become identified as projective representations [67]. The projective part is classified by $H^{2}[G;U(1)]$, as is well known. The induced homomorphism from evaluation on a symmetric product state, $e_{0}^{*}$, just corresponds to sending a projective representation to its classifying cohomology class in $H^{2}[G;U(1)]$, and thus reproduces the group cohomology classification of invertible states in one dimension [71, 66].
@@ V Models
In Sec. V.1 we briefly describe a simple class of examples, building on the anomalous Floquet phases of Refs. [22, 23]. These examples are built from swindles applied to a translation QCA, and give a simple picture illustrating the anomalous edge dynamics of ALT phases. Section V.2 constructs a more experimentally viable implementation of the swindle map, and Sec. V.3 applies it to a very detailed investigation of a one-dimensional SALT phase with two drives and $U(1)$-symmetry, the quasiperiodic energy pump (QP pump) [31, 32, 53]. We include explicit expressions for the Hamiltonian and quantized observables which reveal the anomalous edge dynamics. The model we obtain through the swindle construction is relatively simple, at least in its operator content, paving the way for experiments on this phase.